# Robot Pouring: Identifying Causes of Spillage and Selecting Alternative Action Parameters Using Probabilistic Actual Causation

### v1

@@ Abstract
Considering that causal probability computations cannot tell per se what caused an observed outcome as this may be specific to the given circumstances, in Section 4.1, we analyze the causal probability functions for interventions on each variable to gain insight into their influence on spillage and to emphasize their limitations for action guidance. Subsequently, Section 4.2 presents the results of applying actual causation analysis on spillage trials to select alternative action parameters, where some alternative actions yield lower or higher probabilities of spillage. We show that the analysis based on actual causation can be used in a principled way to select alternative action parameters. Finally, in Section 5, we discuss the implications of the variables’ representation and how the alternative actions suggested by the actual causation analysis would compare to the alternative solutions proposed by a human observer.
@@ 1 Probabilistic Actual Causation and Action-Guiding Explanations
The concept of actual cause refers to the conditions under which a particular event is recognized to be responsible for producing an outcome in a specific scenario or context (Pearl, 2009a). The purpose of determining an actual cause is to find which past actions explain an already observed output (Beckers, 2022). The definition of actual causation proposed by Halpern and Pearl (2005) for deterministic scenarios is one of the most prominent definitions in the literature (Borner, 2023). Fenton-Glynn (2017, 2021) proposed an extension of Halpern and Pearl’s definition that is apt for probabilistic causal scenarios. The Fenton-Glynn’s definition of actual causation is formulated in the framework of probabilistic causal models, where the causal model is represented graphically as a directed acyclic graph (DAG) and the link between each node/variable and its direct causes is modeled probabilistically (Pearl, 2009b). In the DAG formalism, nodes correspond to variables, and directed edges (i.e., arrows) indicate causal influences. Additionally, the $do(x)$ operator represents the operation of setting the value of a variable $X$ to $X=x$ (i.e., the variable $X$ is instantiated to a value $x$), such that $P(y|do(x))$ represents the probability of obtaining the outcome $Y$ that would result from setting $X$ to $x$ by means of an intervention. Fenton-Glynn (2021) proposes the following definition of actual causation [^1] The notation and terminology follow the definition PC1 presented by Fenton-Glynn (2021, p. 72):
@@ 1.1 Using the Actual Causation Inequality to Select Alternative Actions
Fenton-Glynn’s framework of actual causation provides contrastive explanations in a given context (Borner, 2023). By definition, inequality (1) demands that the explanation remains valid when holding fixed all variables in $\mathbf{W}$ and (the subsets) $\mathbf{Z^{\prime}}$ at their actual values; probability raising need not hold for other contexts. In this work, we propose to use the contrastive explanations obtained by applying the actual causation inequality for action guidance. Given that an outcome $Y$ has been observed, the goal is to select an alternative action that will prevent the outcome. Recalling that probability raising entails that $X$ taking the value $X=x$ rather than $X=x^{\prime}$ is the cause of event $Y$, we propose to select $x^{\prime}$ values as alternative actions. In the case illustrated in Figure 1, while selecting an $X$ value in the shaded area much smaller than the actual $X=x$ is likely to change the outcome, selecting $X$ close and above the actual value will likely leave the outcome unchanged. We emphasize that the magnitude of probability raising might drastically differ within the range of contrastive values. This has implications for the suitability of different $x\prime$ values as alternative actions. Consider the alternative actions $X=-5$ and $X=0$ in the schematic example illustrated in Figure 1. Based on the probabilities, it can be assumed that $X=-5$ is a better choice since selecting $X=0$ has a higher chance of leaving the outcome unchanged.
In summary, to check for actual causation using inequality (1), one must decide which variable takes the role of $X$ and, given the graph structure, identify the variables for the sets $\mathbf{W}$ and $\mathbf{Z^{\prime}}$. Thus, the analysis of actual causation can be applied to different variables. In the robot pouring task, if spillage occurs, the aim is to find alternative parameters to repeat the action. In this work, we use the probabilistic actual causation framework to guide the selection of alternative parameters in a principled way. We illustrate the different impacts of the choice of alternative actions with a series of examples.
@@ 2 Related Work
This section presents applications that use causal models and the Halpern-Pearl definition of actual causation (Halpern and Pearl, 2005). We focus on robotics and automatic systems. The reviewed works show that practical applications of the concepts of actual causation have been focused on generating explanations, leaving the potential to guide decision-making processes aside. To the best of our knowledge, the probabilistic actual causation definition of Fenton-Glynn (2017, 2021) has not been used in any practical application.
@@ 3.2 DAG Variables
We also consider the fact that spillage occurs when the amount of marbles exceeds the capacity of the target container. The source containers’ capacity and fullness level determine the poured amount (recall that all the marbles are poured). We represent the difference in volume between the poured amount and the target container’s capacity with the variable relative volume ($RV$). Finally, we represent the outcome of the pouring trial with the variable spillage ($S$), a binary variable that indicates whether or not spillage occurred.
@@ 3.3 Determination of DAG Structure Using Causal Discovery
In order to determine the structure of the DAG, the variables described in Section 3.2 were processed with the PC algorithm (Spirtes and Glymour, 1991), a well-established causal discovery algorithm (Glymour et al., 2019; Nogueira et al., 2022). In general, causal discovery algorithms, also known as structure learning algorithms, perform a systematic analysis of many possible causal structures, typically by testing probabilistic independence and dependence between the variables (Malinsky and Danks, 2017; Glymour et al., 2019; Nogueira et al., 2022). We used the PC implementation available in Tetrad (version 7.6.5-0) [^3] Publicly available at: https://www.ccd.pitt.edu/tools/ , access: 06.12.24, a software toolbox for causal discovery (Ramsey et al., 2018). As a statistical test, we use the Degenerate Gaussian Likelihood Ratio Test (DG-LRT) (Andrews et al., 2019), which has shown good discovery performance on datasets containing continuous and discrete variables (Andrews et al., 2019). We executed the algorithm on 1000 bootstraps of the data to ensure the stability and reliability of the inferred causal relationships (if the results vary widely over the different bootstrap samples, the output of the algorithm is considered unstable) (Glymour et al., 2019). Further details about the PC parameters and bootstrapping results are provided in the supplementary materials. The discovered DAG is shown in Figure 3.
@@ 3.5.1 Causal Probability Expressions
The DAG indicates that $S$ has direct and indirect causes. The path $RC\rightarrow RV\rightarrow S$ shows that $RC$ is an indirect cause of $S$, and its effect is mediated by $RV$. The causal relation between $FU$ and $S$ combines a direct path ($FU\rightarrow S$) and indirect path ($FU\rightarrow RV\rightarrow S$). Finally, $RD$ is a direct cause of S (path $RD\rightarrow S$). Ultimately, the probability of $S$ depends on the interactions between $RV$, $FU$, and $RD$. Additional insight into the factors that yield spillage can be gained by estimating the effects of each variable on $S$. Causal effects are functions of interventional probabilities $P(S|do(X)$ over different values of $X$; the expressions can be obtained using the rules of $do$-calculus with the causaleffect R package (version 1.3.15) (Tikka and Karvanen, 2017). Following the procedure done by Garrido et al. (2021), we use Monte Carlo integration in order to obtain a numerical approximation of the causal expressions of interest.
@@ 4.2 Actual Causes of Spillage and Alternative Actions
According to the considerations presented in Section 3.2 and the DAG discovered from the training data (Figure 3), spillage has three direct causes: $RV$, $FU$, and $RD$. At the beginning of the pouring action, i.e., at the pouring onset, the interaction of $FU$ and $RD$ determines the probability of spillage. Afterward, as the content is poured into the target container, spillage will occur after the poured amount exceeds the container’s capacity; that is, spillage caused by $RV$. Recalling that $RV$ is caused by $FU$ and $RC$, the capacity of the target will be exceeded only if $RC<1$. The smaller the $RC$, the smaller the $FU$ will be necessary to exceed the target container’s capacity. Otherwise, when $RC>1$, spillage is caused only by $FU$ and $RD$. Therefore, we exemplify the usage of the actual causation framework for the analysis of spillage trials with $10.5<RC<$ 1. For simplicity, we focus the analysis on $RD$ and $FU$.
As explained in Section 1, the actual causation inequality compares a reference probability value against the probabilities obtained from contrastive values to check whether or not probability raising holds. As proposed in Section 1.1, we use the contrastive values where probability raising holds to guide the selection of alternative parameters to avoid spillage.
In each example, we start by presenting the frequency of the outcomes (spillage true or false) over 100 replications using the actual trial parameters. Afterward, we perform the actual causation analysis of the $FU$ and $RD$ variables. To compare the suitability of different alternative parameters, we select alternative $FU$ or $RD$ parameters where $P(Y|do(\mathbf{W=w^{*}},X=x^{\prime})$ yields a low probability (0.2) and a chance-level (0.5) probability. We conduct 100 replications using the alternative parameters and present the frequency of the outcomes.
@@ 4.2.1 Example 1
Figure 6(c) shows the reference probability and the probabilities obtained for contrastive $RD$ values obtained from inequality (9). In the area where probability raising holds, the contrastive probabilities show a sharp decrease around $RD\approx 0.8$ and a low probability for $RD>0.9$. We select the alternative values $RD=0.87$ (chance-level probability) and $RD=0.89$ (low probability $\approx 0.2$). It is important to note that changing $RV$ and keeping $RC$ constant produce a target container of smaller height. The results from 100 replications are shown in Figure 6(f) and Figure 6(h). The replications with the alternative parameters show that $RD=0.89$ is better suited to avoid spillage. It is also interesting to note that due to the non-linearity of the probabilities, small changes in $RD$ yield a large effect on the probability of spillage.

### v3

@@ 1 Introduction
Considering that causal probability computations cannot tell per se what caused an observed outcome as this may be specific to the given circumstances, in Section 5.1, we analyze the causal probability functions for interventions on each variable to gain insight into their influence on spillage and to emphasize their limitations for action guidance. Subsequently, in Sections 5.2 and 5.3 we demonstrate that the analysis based on actual causation can be used in a principled way to select alternative action parameters. Firstly, in Section 5.2 we present four detailed examples of applying actual causation analysis on spillage trials to select alternative action parameters. The examples illustrate how different alternative actions yield lower or higher probabilities of spillage. Secondly, in Section 5.3 we evaluate the likelihood of finding an alternative value and the pouring success rates obtained when running the spillage trials using alternative values. The evaluation was conducted using a test dataset acquired in simulation. Overall, Sections 5.2 and 5.3 provide empirical evidence of the practical usefulness of the actual causation approach to identify alternative parameters to prevent spillage.
Finally, in Section 6, we discuss the implications of the variables’ representation and how the alternative actions suggested by the actual causation analysis would compare to the alternative solutions proposed by a human observer.
In summary, our contributions are:
  • We report a complete analysis of probabilistic actual causation on a practical problem with all the necessary steps for its implementation (analysis of the task, variable definition, causal-graph structure, and estimation of causal probabilities using neural networks).
  • We show how to use the explanations obtained with the analysis of probabilistic actual causation for action guidance in a practical use case. By doing this, we go beyond the basic diagnostic (attribution of actual cause) toward action guidance.
  • We evaluate the capability of the actual causation approach to automatically identify alternative parameters of different variables.
  • We evaluate the pouring success rates obtained when the alternative parameters are used to prevent spillage.
@@ 2 Probabilistic Actual Causation and Action-Guiding Explanations
The concept of actual cause refers to the conditions under which a particular event is recognized to be responsible for producing an outcome in a specific scenario or context (Pearl, 2009a). The purpose of determining an actual cause is to find which past actions explain an already observed output (Beckers, 2022). The definition of actual causation proposed by Halpern and Pearl (2005) for deterministic scenarios is one of the most prominent definitions in the literature (Borner, 2023). Fenton-Glynn (2017, 2021) proposed an extension of Halpern and Pearl’s definition that is apt for probabilistic causal scenarios. The Fenton-Glynn’s definition of actual causation is formulated in the framework of probabilistic causal models, where the causal model is a causal Bayesian network represented graphically as a directed acyclic graph (DAG) and the link between each node/variable and its direct causes is modeled probabilistically (Pearl, 2009b). In the DAG formalism, nodes correspond to variables, and directed edges (i.e., arrows) indicate causal influences. Additionally, the $do(x)$ operator represents the operation of setting the value of a variable $X$ to $X=x$ (i.e., the variable $X$ is instantiated to a value $x$), such that $P(y|do(x))$ represents the probability of obtaining the outcome $Y$ that would result from setting $X$ to $x$ by means of an intervention. Fenton-Glynn (2021) proposes the following definition of actual causation [^1] The notation and terminology follow the definition PC1 presented by Fenton-Glynn (2021, p. 72):
@@ 2.1 Using the Actual Causation Inequality to Select Alternative Actions
Fenton-Glynn’s framework of actual causation provides contrastive explanations in a given context (Borner, 2023). By definition, inequality (1) demands that the explanation remains valid when holding fixed all variables in $\mathbf{W}$ and (the subsets) $\mathbf{Z^{\prime}}$ at their actual values; probability raising need not hold for other contexts. In this work, we propose to use the contrastive explanations obtained by applying the actual causation inequality for action guidance. Given that an outcome $Y$ has been observed, the goal is to select an alternative action that will prevent the outcome. Recalling that probability raising entails that $X$ taking the value $X=x$ rather than $X=x^{\prime}$ is the cause of event $Y$, we propose to select $x^{\prime}$ values as alternative actions. In the case illustrated in Figure 1, while selecting an $X$ value in the shaded area much smaller than the actual $X=x$ is likely to change the outcome, selecting $X$ close and above the actual value will likely leave the outcome unchanged. We emphasize that the magnitude of probability raising might drastically differ within the range of contrastive values. This has implications for the suitability of different $x\prime$ values as alternative actions. Consider the alternative actions $X=-5$ and $X=0$ in the schematic example illustrated in Figure 1. Based on the probabilities, it can be assumed that $X=-5$ is a better choice since selecting $X=0$ has a higher chance of leaving the outcome unchanged. Under these considerations, the selection of an alternative parameter can be based on a pre-defined probability threshold.
An alternative parameter can be found automatically with the following steps:
  1. 1.
    Identify the range of contrastive values where probability raising holds using inequality (1).
  2. 2.
    Within these values, select the subset of values with a probability below a pre-defined probability threshold.
  3. 3.
    From this second subset, select the closest value to the current parameter.
In the last step we use a distance criterion. However, other application-dependent criteria can be used (e.g., the cost or availability of different alternatives) to select an alternative parameter. The crucial aspect of the automatic search is that the probability threshold is selected such that it becomes very likely to change the outcome.
In summary, to check for actual causation using inequality (1), one must decide which variable takes the role of $X$ and, given the graph structure, identify the variables for the sets $\mathbf{W}$ and $\mathbf{Z^{\prime}}$. Thus, the analysis of actual causation can be applied to different variables. In the robot pouring task, if spillage occurs, the aim is to find alternative parameters to repeat the action. In this work, we use the probabilistic actual causation framework to guide the selection of alternative parameters in a principled way. In Section 5.2, we illustrate in a series of detailed examples the impact of using different probability thresholds (i.e., $0.2$ for low spillage probability and and $0.5$ for chance-level probability) on the identified alternative parameters. Subsequently, in Section 5.3, we evaluate the pouring success rates obtained using the alternative parameters identified using a $0.1$ probability threshold.
@@ 3 Related Work
This section presents robotics applications that use causal methods related to our work (i.e., actual causation, causal Bayesian networks, and causal discovery). We start by presenting applications that use the Halpern-Pearl definition of actual causality (Halpern and Pearl, 2005). Subsequently, we review works that use causal discovery to learn the structure of causal Bayesian networks to model robotic tasks in different contexts. Finally, we review the method proposed by Diehl and Ramirez-Amaro (2023) to predict and prevent failures using causal-based contrastive explanations, highlighting the similarities and differences to our approach.
@@
The reviewed works show that practical applications of the concepts of actual causation have focused on generating explanations, leaving the potential to guide decision-making processes aside. To the best of our knowledge, the probabilistic actual causation definition of Fenton-Glynn (2017, 2021) has not been used in any practical application.
Causal methods have been applied to model generic robot tasks such as pushing, pick-and-place, and stacking (Ahmed et al., 2020; Brawer et al., 2020; Huang et al., 2023; Diehl and Ramirez-Amaro, 2023) and context-specific ones such as human-robot interaction Castri et al. (2022) and household tasks (Li et al., 2020). In the context of tool affordance learning, causal discovery has been used to identify the effect of push and pull actions on an object’s final position in a real setup (Brawer et al., 2020). To reduce the sim-to-real gap in robot-object trajectories, a custom causal discovery algorithm was used to optimize the simulated physical parameters (Huang et al., 2023). Additionally, recent work has used simulations to explore large combinatoric spaces of task parameters and causal methods to model task outcomes (Ahmed et al., 2020; Huang et al., 2023; Diehl and Ramirez-Amaro, 2023).
Similar to our work, causal methods have been applied to find causal-based contrastive explanations for task failures (Diehl and Ramirez-Amaro, 2022). In subsequent work, this method has been extended to predict and prevent task failures by finding corrective parameters (Diehl and Ramirez-Amaro, 2023). Examples of contrastive explanations are provided for a cube stacking task and for dropping spheres into different containers (bowls, plates, and glasses) (Diehl and Ramirez-Amaro, 2022), and the method for prediction and prevention has been applied to a cube stacking task (stacking one cube and stacking three cubes) (Diehl and Ramirez-Amaro, 2023).
In particular, the method proposed by Diehl and Ramirez-Amaro (2023) has methodological similarities to our approach. Their method uses a causal Bayesian network to predict errors and probabilities of success to find a corrective action. Simulated data are used to learn the causal Bayesian network’s structure and estimate its joint probability distribution. When an action is predicted to fail, a search is conducted in a discretized parameter space to find the parametrization that needs the least interval changes to achieve a successful execution based on the predicted success probability (Diehl and Ramirez-Amaro, 2023). For example, searching for close parametrizations in the cube-stacking task means that starting from a robot hand position that will likely fail leads to finding the closest position that will likely succeed (Diehl and Ramirez-Amaro, 2023).
The search criterion is based on contrastive explanations (Diehl and Ramirez-Amaro, 2022), which compare the variable parametrization of the failed action and the closest parametrization that exceeds a success probability threshold ($\epsilon$(Diehl and Ramirez-Amaro, 2023). The success probabilities are retrieved from the factorized form of the joint probability distribution. The search is conducted within a tree that contains a complete parametrization of the parent variables of the outcome variable. Since the search is conducted within a tree structure, the time to find an alternative parametrization depends on the success probability threshold, the number of parent variables, and the number of discrete intervals of the variables (Diehl and Ramirez-Amaro, 2023).
In Table 1, we summarize the similarities and differences between the method proposed by Diehl and Ramirez-Amaro (2023) and our approach. Their method automatically finds the corrective parameters, which might involve changing the value of one or more variables. In contrast, our approach is restricted to searching for corrective parameters in a single variable. Their method identifies the closest parametrization that is predicted to succeed. On the other hand, our approach identifies a range of values that explain the outcome according to the probability-raising criterion. Within this range of values, a specific value likely to yield success can be selected using a probability criterion (see explanation in Section 2.1). Similarly to Diehl and Ramirez-Amaro (2023), a success probability threshold can be used as a selection criterion.
Table 1: Comparison of methods
Similarities Differences
Model Training data Type of explanation Selection of corrective action Input Parameter space Corrected variables Contrasted probabilities
Diehl and Ramirez-Amaro (2023) Causal Bayesian network (discrete variables) Simulation Contrastive Based on a probability threshold Current values of all variables (i.e., parametrization) and probability of the outcome variable Discrete The corrective parametrization might involve changing one or more variables Conditional probabilities from the factorized joint probability distribution
Ours Causal Bayesian network (continuous and discrete variables) Simulation Contrastive Based on a probability threshold Current values of all variables (i.e., parametrization), causal probability (do-operator) of the outcome variable, and variable to correct Continuous The corrective parametrization changes one variable Causal probabilities (do-operator)
@@ 4.2 DAG Variables
We also consider the fact that spillage occurs when the amount of marbles exceeds the capacity of the target container. The source containers’ capacity and fullness level determine the poured amount (recall that all the marbles are poured). We represent the difference in volume between the poured amount and the target container’s capacity with the variable relative volume ($RV$). Finally, we represent the outcome of the pouring trial with the variable spillage ($S$), a binary variable that indicates whether or not spillage occurred. $S$ is labeled as $true$ irrespective of the number of spilled marbles (i.e., spilling one or twenty marbles yields $S=true$).
@@ 4.3 Determination of DAG Structure Using Causal Discovery
In order to determine the structure of the DAG, the variables described in Section 4.2 were processed with the PC algorithm (Spirtes and Glymour, 1991), a well-established causal discovery algorithm (Glymour et al., 2019; Nogueira et al., 2022). In general, causal discovery algorithms, also known as structure learning algorithms, perform a systematic analysis of many possible causal structures, typically by testing probabilistic independence and dependence between the variables (Malinsky and Danks, 2017; Glymour et al., 2019; Nogueira et al., 2022). We used the PC implementation available in Tetrad (version 7.6.5-0) [^3] Publicly available at: https://www.ccd.pitt.edu/tools/ , access: 06.12.24, a software toolbox for causal discovery (Ramsey et al., 2018). As a statistical test, we use the Degenerate Gaussian Likelihood Ratio Test (DG-LRT) (Andrews et al., 2019), which has shown good discovery performance on datasets containing continuous and discrete variables (Andrews et al., 2019). We executed the algorithm on 1000 bootstraps of the data to ensure the stability and reliability of the inferred causal relationships (if the results vary widely over the different bootstrap samples, the output of the algorithm is considered unstable) (Glymour et al., 2019). Further details about the PC parameters, bootstrapping results, and modeling assumptions are provided in the supplementary materials. The discovered DAG is shown in Figure 3.
@@ 4.5.1 Causal Probability Expressions
The DAG indicates that $S$ has direct and indirect causes. The path $RC\rightarrow RV\rightarrow S$ shows that $RC$ is an indirect cause of $S$, and its effect is mediated by $RV$. The causal relation between $FU$ and $S$ combines a direct path ($FU\rightarrow S$) and indirect path ($FU\rightarrow RV\rightarrow S$). Finally, $RD$ is a direct cause of S (path $RD\rightarrow S$). Ultimately, the probability of $S$ depends on the interactions between $RV$, $FU$, and $RD$. Additional insight into the factors that yield spillage can be gained by estimating the effects of each variable on $S$. Causal effects are functions of interventional probabilities $P(S|do(X))$ over different values of $X$; the expressions can be obtained using the rules of $do$-calculus with the causaleffect R package (version 1.3.15) (Tikka and Karvanen, 2017). Following the procedure done by Garrido et al. (2021), we use Monte Carlo integration in order to obtain a numerical approximation of the causal expressions of interest.
@@ 5.2 Actual Causes of Spillage and Alternative Actions
As explained in Section 2, the actual causation inequality compares a reference probability value against the probabilities obtained from contrastive values to check whether or not probability raising holds. As proposed in Section 2.1, we use the contrastive values where probability raising holds to guide the selection of alternative parameters to avoid spillage. In this section, we present the results obtained by using the automatically-selected alternative values to achieve either a chance-level or a low probability of spillage. In the following, we explain the rationale behind the selected spillage trial examples.
According to the considerations presented in Section 4.2 and the DAG discovered from the training data (Figure 3), spillage has three direct causes: $RV$, $FU$, and $RD$. At the beginning of the pouring action, i.e., at the pouring onset, the interaction of $FU$ and $RD$ determines the probability of spillage. Afterward, as the content is poured into the target container, spillage will occur after the poured amount exceeds the container’s capacity; that is, spillage caused by $RV$. Recalling that $RV$ is caused by $FU$ and $RC$, the capacity of the target will be exceeded only if $RC<1$. The smaller the $RC$, the smaller the $FU$ will be necessary to exceed the target container’s capacity. Otherwise, when $RC>1$, spillage is caused only by $FU$ and $RD$. Therefore, we exemplify the usage of the actual causation framework for the analysis of spillage trials with $10.5<RC<$ 1. For simplicity, we focus the analysis on $RD$ and $FU$, which have a direct effect on $S$. We do not consider $RC$ because it doesn’t have a direct effect on $S$ (the effect of $RC$ is mediated by $RV$, which also depends on $FU$).
In each example, we start by presenting the frequency of the outcomes (spillage true or false) over 100 replications using the actual trial parameters. Afterward, we perform the actual causation analysis of the $FU$ and $RD$ variables. To compare the suitability of different alternative parameters, we select alternative $FU$ or $RD$ parameters where $P(Y|do(\mathbf{W=w^{*}},X=x^{\prime})$ yields a low probability (0.2) and a chance-level (0.5) probability. The alternative parameters are identified automatically following the steps described in Section 2.1. We conduct 100 replications using the alternative parameters and present the frequency of the outcomes.
@@ 5.2.1 Example 1
Figure 6(c) shows the reference probability and the probabilities obtained for contrastive $RD$ values obtained from inequality (9). In the area where probability raising holds, the contrastive probabilities show a sharp decrease around $RD\approx 0.8$ and a low probability for $RD>0.9$. We select the alternative values $RD=0.87$ (chance-level probability) and $RD=0.89$ (low probability $\approx 0.2$), as shown in Figures 6(e) and 6(g), respectively. It is important to note that changing $RV$ and keeping $RC$ constant produce a target container of smaller height. The results from 100 replications are shown in Figure 6(f) and Figure 6(h). The replications with the alternative parameters show that $RD=0.89$ is better suited to avoid spillage. It is also interesting to note that due to the non-linearity of the probabilities, small changes in $RD$ yield a large effect on the probability of spillage.